# Universal consistency of the $k$-NN rule in metric spaces and Nagata dimension. III

### v1

@@ 5.1.1. Step $m_{i}\to n_{i}$
By the Chernoff bound, the probability of the empirical measure of the singleton $\{x^{t}\}$ deviating from its expectation, $\mu\{x^{t}\}$, by more than half its value is limited by
@@
Choose any value $n=n_{i}$ satisfying the last inequality and such that eq. (4) holds for $n\geq n_{i}$. If now $n\geq n_{i}$, then, with confidence $1-\delta_{i}$, the empirical measure of every point $x^{t}$, $\lvert t\rvert=i$ will be at least $(1/2)\mu\{x^{t}\}>k_{n_{i}}/n_{i}$. Consequently, each point $x^{t}$, $\lvert t\rvert=i$ will appear at least $k_{n}$ times among the members of a random i.i.d. $n$-sample following the law $\mu$.
@@ 5.1.2. Step $n_{i}\to m_{i+1}$
Now we will choose $m_{i+1}$. We want this number to be so large that the proportion of the number of balls $B_{{\epsilon}^{t}}(y^{tj})$, $\lvert t\rvert=i$, $j=1,2,\ldots,m_{i+1}$, containing at least $k_{n_{i}}/2$ elements of a random i.i.d. $n_{i}$-sample following the law $\mu$ is less than $\delta_{i}$. The maximal possible number of such balls is limited by $2n_{i}/k_{n_{i}}$. So it is enough to require that the ratio of $2n_{i}/k_{n_{i}}$ to the total number of points $y^{t}$ with $\lvert t\rvert=i+1$ (that is, the number of balls under consideration) is less than $\delta_{i}$:
$\frac{2n_{i}}{k_{n_{i}}}<\delta_{i}\prod_{j=0}^{i+1}m_{j}=\delta_{i}\left(\prod_{j=0}^{i}m_{j}\right)m_{i+1}.$
This means selecting $m_{i+1}$ with
@@
Notice in passing that as $n_{i}/\prod_{j=0}^{i}m_{j}\geq k_{n_{i}}$ by eq. (4), the r.h.s. of the above expression is no less than $2/\delta_{i}$, and so $m_{i}\to\infty$. (This assures that the measure $\mu_{0}$ is diffuse, see the end of Subs. 4.3.2, though for that it would suffice that each $m_{i}\geq 2$.)
5.2. Absence of weak consistency of the $k$-NN rule under $(\mu,\eta)$
5.2.1. Predictor for $x^{t}$
Each point $x^{t}$ is an atom, labelled deterministically $1$. Under such conditions, it is well known and easily proved that the label predicted by the $k$-NN classifier in the limit $n\to\infty$ converges on the set of atoms in probability to $1$.
It remains to analyse the situation with $x\in K$ that is not of the form $x^{t}$ and thus will $\mu$-almost surely be labelled 0. We will show that the label predicted by the $k$-NN classifier does not converge in probability to zero.
5.2.2. $x\in K\setminus\{x^{t}\colon t\in T\}$
@@
Going back to Subs. 4.3, we notice that, since each point $y^{t}$ has zero measure, $\mu$-almost all points $x$ of the set $K\setminus\{x^{t}\colon t\in T\}$ (in fact, all except a finite number) are contained in the set $\cup_{\lvert t\rvert=i}\cup_{j=1}^{m_{i+1}}B_{{\epsilon}^{t}}(y^{tj})$. Since the balls are pairwise disjoint, for $\mu$-almost all $x$ there are unique $t\in T$ with $\lvert t\rvert=i$ and $j=1,2,\ldots,m_{i+1}$ such that $x\in B_{{\epsilon}^{t}}(y^{tj})$.
Now, $x^{t}$ is strictly closer to $x$ than any point of any other ball of the same form, $B_{{\epsilon}^{t}}(y^{ts})$, $s\neq j$ (property 4.2(1)). If we form a closed ball around $x$ of radius $d(x,x^{t})$, the intersection of this ball with $K$ will be the same as the intersection of the ball $B_{{\epsilon}^{t}}(y^{tj})$ with $K$ plus $\{x^{t}\}$.
Hence, with probability $1-\delta_{i}$, conditionally on $X\notin\{x^{t}\colon t\in T\}$, the open ball around a random element $X\sim\mu$ having the same radius as the distance to the nearest $x^{t}$, $\lvert t\rvert=i$, contains strictly fewer than $k_{n_{i}}/2$ elements of the random i.i.d. $n_{i}$-sample following the law $\mu$ (Subs. 5.1.2).
At the same time, with probability $1-\delta_{i}$, each of the points $x^{t}$, $\lvert t\rvert=i$ appears at least $k_{n_{i}}$ times among the elements of the random i.i.d. $n_{i}$-sample following the law $\mu$ (Subs. 5.1).
@@
This means that, with probability $1-2\delta_{i}$, a random element $X\notin\{x^{t}\colon t\in T\}$ will get labelled $1$ by the $k$-NN classifier at the step $(n_{i},k_{n_{i}})$.
5.2.3. As $i\to\infty$
@@
As $\sum\delta_{i}<\infty$, by Borel–Cantelli, the subsequence of $k$-NN classifiers corresponding to $(n_{i},k_{n_{i}})$ converges to the identical value $1$ over $\Omega\setminus\{x^{t}\colon t\in T\}$ almost surely, in particular, in probability. Combining this with the observation in Subs. 5.2.1, we conclude: the subsequence of $k$-NN classifiers corresponding to $(n_{i},k_{n_{i}})$ converges to the identical value $1$ over all of $\Omega$ in probability.
@@
This implies that the sequence of $k$-NN classifiers corresponding to the original sequence $(n,k_{n})$ does not converge in probability to the Bayes classifier, $\eta$, and so the $k$-NN rule is not weakly consistent under $(\mu,\eta)$.
6. Weak Lebesgue–Besicovitch property is not sufficient for the consistency of the $k$-NN classifier
Here we aim to correct a statement made in the previous article [11], namely, Corollary 3.18 in Section 3.
In that section, we reminded the weaker notion of dimension of a metric space in the sense of de Groot [5], which is always less than or equal to the Nagata dimension. According to a result of Assouad and Quentin de Gromard ([1], Prop. 3.3.1(b)+Prop. 3.1), a complete separable metric space having finite de Groot dimension satisies the weak Lebesgue–Besicovitch differentiation property for every locally finite Borel measure. A classical example of a metric space of finite de Groot dimension (moreover, having a doubling property) is the Heisenberg group $\mathbb{H}$}, equipped with the Cygan-Korányi metric [6]. At the same time, as shown, independently, by Korányi and Reimann ([9], p. 17) and Sawyer and Wheeden ([19], Lemma 4.4, p. 863), the Nagata dimension of the Heisenberg group $\mathbb{H}$} is infinite, and in fact their argument shows that $\mathbb{H}$} is not sigma-finite dimensional in the sense of Nagata ([11], Corol. 3.20). For a detailed presentation of all of this, with proofs and references, we refer the reader to Sect. 3 in [11].
Now, in [11], we combined the above with the main result as stated in [3] to claim that the $k$-NN classifier in the Heisenberg group is universally consistent. The main result of the present paper, Thm. 1, shows this not to be true.
Since the Heisenberg group $\mathbb{H}$}, equipped with the Cygan-Korányi metric, is not sigma-finite dimensional in the sense of Nagata by [9, 19], the $k$-NN classifier is not weakly consistent in it – even if, according to [1], the group $\mathbb{H}$} with the Cygan-Korányi metric satisfies the weak Lebesgue–Besicovitch differentiation property for every locally finite measure. Modulo our measure construction in Section 4, this gives an example of a complete separable metric space equipped with a probability measure that satisfies the weak Lebesgue–Besicovitch differentiation property but in which the $k$-NN classifier is not consistent.

### v2

@@ 4.2.1. Step $m_{i}\to n_{i}$
By the Chernoff bound, the probability for the empirical measure of the singleton $\{x^{t}\}$ to be smaller than half its expectation, $\mu\{x^{t}\}$, is limited by
@@
Choose any value $n=n_{i}$ satisfying the above inequality and such that eq. (4) holds. If now $n=n_{i}$, then, with confidence $1-\delta_{i}$, the empirical measure of every point $x^{t}$, $\lvert t\rvert=i$ supported on a random $n_{i}$-sample will be at least $(1/2)\mu\{x^{t}\}>k_{n_{i}}/n_{i}$. In other words, with confidence $1-\delta_{i}$, each point $x^{t}$, $\lvert t\rvert=i$ will appear at least $k_{n_{i}}$ times among the members of a random i.i.d. $n_{i}$-sample following the law $\mu$.
@@ 4.2.2. Step $n_{i}\to m_{i+1}$
Now we will choose $m_{i+1}$. We want this number to be so large that no more than a fraction of $\delta_{i}$ among the balls of the form $B_{{\epsilon}^{t}}(y^{tj})$, $\lvert t\rvert=i$, $j=1,2,\ldots,m_{i+1}$ contain at least $k_{n_{i}}/2$ elements of a random i.i.d. $n_{i}$-sample following the law $\mu$. The number of balls containing $\geq k_{n_{i}}/2$ elements of a $n_{i}$-sample is limited by $2n_{i}/k_{n_{i}}$. So it is enough to require
$\frac{2n_{i}}{k_{n_{i}}}<\delta_{i}\prod_{j=0}^{i+1}m_{j}=\delta_{i}\left(\prod_{j=0}^{i}m_{j}\right)m_{i+1},$
that is,
@@
Notice in passing that as $n_{i}/\prod_{j=0}^{i}m_{j}\geq k_{n_{i}}$ by eq. (4), the r.h.s. of the above expression is no less than $2/\delta_{i}$, and so $m_{i}\to\infty$. (This assures that the measure $\mu_{0}$ is diffuse, see the end of Subsection 4.1.7, though for that it would already suffice that each $m_{i}\geq 2$.)
4.3. Absence of consistency of the $k$-NN rule under $(\mu,\eta)$
4.3.1. Predictor for atoms
Each point $x^{t}$ is an atom, labelled deterministically $1$. Under such conditions, it is well known and easily proved that on the set of atoms the $k$-NN classifier in the limit $n\to\infty$ will converge in probability to the identical value $1$, whether $k$ goes to infinity or is even fixed.
It remains to analyse the situation with $x\in K$ that is not of the form $x^{t}$ and thus will $\mu$-almost surely be labelled 0. We will show that on such points, the label predicted by the $k$-NN classifier does not converge in probability to zero.
4.3.2. Pedictor for non-atomic points
@@
Going back to Subsection 4.1.5, we notice that, since each point $y^{t}$ has zero measure, $\mu$-almost all elements $x$ of the non-atomic part of $K$, that is, of $K\setminus\{x^{t}\colon t\in T\}$ (in fact, all except possibly a finite number) are contained in the set $\cup_{\lvert t\rvert=i}\cup_{j=1}^{m_{i+1}}B_{{\epsilon}^{t}}(y^{tj})$. Since the balls are pairwise disjoint, for $\mu$-almost every non-atomic $x$ there is a unique $t\in T$ with $\lvert t\rvert=i$ and $j=1,2,\ldots,m_{i+1}$ such that $x\in B_{{\epsilon}^{t}}(y^{tj})$.
Now, $x^{t}$ is strictly closer to $x$ than any point of any other ball of the same form, $B_{{\epsilon}^{t}}(y^{ts})$, $s\neq j$ (property 4.1.2(1)). If we form a closed ball around $x$ of radius $d(x,x^{t})$, the intersection of this ball with $K$ will be the same as the intersection of the ball $B_{{\epsilon}^{t}}(y^{tj})$ with $K$ plus $\{x^{t}\}$.
Hence, with probability $1-\delta_{i}$, conditionally on $X\notin\{x^{t}\colon t\in T\}$, the open ball around a random element $X\sim\mu$ having the same radius as the distance to the nearest $x^{t}$, $\lvert t\rvert=i$, contains strictly fewer than $k_{n_{i}}/2$ elements of the random i.i.d. $n_{i}$-sample following the law $\mu$ (Subsection 4.2.2).
At the same time, with probability $1-\delta_{i}$, each of the points $x^{t}$, $\lvert t\rvert=i$ appears at least $k_{n_{i}}$ times among the elements of the random i.i.d. $n_{i}$-sample following the law $\mu$ (Subs. 4.2.1).
@@
This means that, with probability $1-2\delta_{i}$, a random element $X\in K\setminus\{x^{t}\colon t\in T\}$ will get labelled $1$ by the $k$-NN classifier at the step $(n_{i},k_{n_{i}})$.
4.3.3. $i\to\infty$
@@
As $\sum\delta_{i}<\infty$, by Borel–Cantelli, the subsequence of $k$-NN classifiers corresponding to $(n_{i},k_{n_{i}})$ converges to the identical value $1$ over $\Omega\setminus\{x^{t}\colon t\in T\}$ almost surely, in particular, in probability. Combining this with the observation in Subs. 4.3.1, we conclude: the subsequence of $k$-NN classifiers corresponding to the values $n=n_{i}$, $k=k_{n_{i}}$ converges to the identical value $1$ over all of $\Omega$ in probability as $i\to\infty$.
@@
This implies that the sequence of $k$-NN classifiers corresponding to the original sequence $(n,k_{n})$ does not converge in probability to the Bayes classifier, $\eta$, and so the $k$-NN rule is not consistent under $(\mu,\eta)$.
5. The weak Lebesgue–Besicovitch property is insufficient for the $k$-NN classifier consistency
Here is another corollary obtained by combining our main result (Theorem 1) with known constructions from real analysis in metric spaces.
{crllr}
There exists a complete separable metric space equipped with a Borel probability measure $\mu$, satisfying the weak Lebesgue–Besicovitch property for every $L^{1}(\mu)$-function, yet in which the $k$-NN classification rule is not universally consistent for any sequence of values of $k_{n}\to\infty$, $k_{n}/n\to 0$.
A source of metric spaces in which every Borel probability measure satisfies the weak Lebesgue–Besicovitch property is given by the following concept, which, just like the Nagata dimension, has originated in the topological dimension theory, but is weaker (less restrictive).
{dfntn}
[[11]; [3], 3.5, 3.7] Let $\delta\in{\mathbb{N}}$. A metric space $\Omega$ has de Groot dimension $\leq\delta$ (on the scale $s=+\infty$) if every finite family of closed balls having the same radii admits a subfamily covering all the centres of the original balls and having multiplicity $\leq\delta+1$. The de Groot dimension of $\Omega$ is the smallest $\delta$ with this property, if it exists, and $\infty$ otherwise. We denote it $\dim_{dG}(\Omega)$.
Comparing this concept with Definition 3, we conclude that de Groot dimension is bounded by the Nagata dimension: $\dim_{dG}(\Omega)\leq\dim_{Nag}(\Omega)$ (on the scale $\infty$, the only one we will need). The distinguishing examples between the two dimensions are easy to construct. For instance, the convergent sequence $2^{-n}e_{n}$, $n=1,2,\ldots$ together with the limit 0 in the Hilbert space $\ell^{2}$ (where $e_{n}$ are standard orthonormal basic vectors) has de Groot dimension $2$ yet its Nagata dimension is infinite (see [21], Example 3.12).
{thrm}
[Assouad and Quentin de Gromard, [3], Prop. 3.3.1(b)+Prop. 3.1] Let a complete separable metric space $\Omega$ have finite de Groot dimension. Then for every locally finite Borel measure $\mu$ on $\Omega$, every $L^{1}(\mu)$-function $f\colon\Omega\to{\mathbb{R}}$ satisfies the weak Lebesgue–Besicovitch differentiation property:
$\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}f(x^{\prime})\,d\mu(x^{\prime})\to f(x)~~\mbox{when $r\to 0$}$ (5)
in measure.
Thus, in order to establish Corollary 5, it is enough, in view of Theorem 1, to give examples of complete separable metric spaces having finite de Groot dimension, which are at the same time not sigma-finite dimensional in the sense of Nagata.
While it may not be so easy to verify the finiteness of de Groot dimension directly, in view of its combinatorial flavour, the following is an important subclass of metric spaces having finite de Groot dimension. Recall that a metric space $(X,d)$ has the doubling property if there is a constant $C\geq 1$ such that for every $r>0$, every closed ball of radius $2r$ is covered by $\leq C$ closed balls of radii $r$.
{prpstn}
A metric space $\Omega$ having the doubling property with constant $C$ has finite de Groot dimension with $\dim_{dG}(\Omega)\leq C-1$.
Proof.
Let $\bar{B}_{r}(x_{i})$, $i=1,2,\ldots,n$ be a finite family of closed balls all having the same radius $r>0$. Select a maximal set of indices, $J\subseteq[n]$, such that the centres $x_{i},x_{j}$ for $\in J, iji\neq j$ are at a distance >r>r between themselves. Clearly, the balls $\bar{B}_{r}(x_{i})$, $i\in J$ still contain all the original centres, by the maximality of $J$. Let now $x\in\Omega$ be any. Cover the ball $\bar{B}_{r}(x)$ with balls $\bar{B}_{r/2}(y_{k})$, $k=1,2,\ldots,C$. Each ball of radius $r/2$ can only contain at most one element $x_{i}$, $i\in J$. This means no more than $C$ balls of the form $\bar{B}_{r}(x_{i})$, $i\in J$ contain $x$, and so the multiplicity of the subfamily $\bar{B}_{r}(x_{i})$, $i\in J$ is $\leq C$. ∎
{xmpl}
[Käenmäki, Rajala and Suomala, Example 5.6 in [16]] There exists a metric $\rho$ on the Cantor space $C$ (realized as a product of a sequence of finite spaces) having the doubling property and yet such that with regard to $\rho$ and a suitable product probability measure on $C$ the Lebesgue–Besicovitch differentiation property fails.
In view of the theorem of Preiss–Assouad–Quentin de Gromard (Subsection 2.2), the metric space $\left(C,\rho\right)$ is not sigma-finite dimensional in the sense of Nagata. At the same time, by Proposition 5, this metric space has finite de Groot dimension. This, in view of Theorem 1, gives perhaps the simplest example with a desired combination of properties to validade Corollary 5.
The second example is classical.
{xmpl}
[Heisenberg group] The Heisenberg group $\mathbb{H}$} is the Euclidean space ${\mathbb{R}}^{3}$ equipped with the following group multiplication:
$(x,y,z)\cdot(x^{\prime},y^{\prime},z^{\prime})=\left(x+x^{\prime},y+y^{\prime},z+z^{\prime}-2xy^{\prime}+2yx^{\prime}\right).$ (6)
The identity of this group is the zero element $(0,0,0)$ of ${\mathbb{R}}^{3}$, and the inverse of $(x,y,z)$ is the same as the usual additive inverse (x,y,z)(-x,-y,-z).
The formula
$\left|(x,y,z)\right|_{{\mathbb{H}}}=\left((x^{2}+y^{2})^{2}+z^{2}\right)^{1/4}$
defines a group norm on $\mathbb{H}$}, in the sense that $\lvert p^{-1}\rvert_{{\mathbb{H}}}=\lvert p\rvert_{{\mathbb{H}}}$ and $\lvert p\cdot q\rvert_{{\mathbb{H}}}\leq\lvert p\rvert_{{\mathbb{H}}}+\lvert q\rvert_{{\mathbb{H}}}$. The latter inequality is not quite trivial to verify, and follows from a more general inequality of Cygan [12]. As a consequence, the formula
$d(p,q)=\lvert p^{-1}\cdot q\rvert_{{\mathbb{H}}},~~p,q\in{\mathbb{H}}$
defines a metric on the group $\mathbb{H}$}, which is easily seen to be left-invariant and compatible with the Euclidean topology. This metric is known as either (Cygan-)Korányi, or else Carnot distance. See e.g. [22] for more.
The doubling property for the above metric immediately follows from the fact it is not only left-invariant, but also homogeneous: for $t>0$, the transformation $(x,y,z)\mapsto(tx,ty,t^{2}z)$ increases the distance between any pair of points by the factor of $t$. In this way, it is enough to cover the closed ball of radius one around identity with finitely many balls of radius $1/2$, and this holds because of local compactness. We conclude that the Heisenberg group has finite de Groot dimension.
At the same time, it was shown, independently, by Korányi and Reimann ([18], p. 17) and Sawyer and Wheeden ([29], Lemma 4.4, p. 863), that the Nagata dimension of the Heisenberg group $\mathbb{H}$} is infinite, and in fact their argument shows more: that $\mathbb{H}$} is not sigma-finite dimensional in the sense of Nagata (see [21], Corol. 3.20).
Thus, the Heisenberg group gives another way to establish Corollary 5 of Theorem 1.
{rmrk}
We use this opportunity to correct a wrong statement made in the previous article [21], namely, Corollary 3.18, where we erroneously claimed that the $k$-NN classifier is universally consistent in the Heisenberg group. The main result of the present paper, Theorem 1, implies otherwise.
6. An equivalent metric on $\mathbb{R}$} under which the $k$-NN classifier fails
Recall that two metrics, $d$ and $\rho$, on the same set $X$ are uniformly equivalent if for every ${\epsilon}>0$ there is $\delta>0$ so that
$\forall x,y\in X,~~d(x,y)<\delta\Rightarrow\rho(x,y)<{\epsilon},~~\rho(x,y)<\delta\Rightarrow d(x,y)<{\epsilon}.$
In particular, such $d$ and $\rho$ are equivalent, that is, induce the same topology. Two uniformly equivalent metrics share the same Cauchy sequences, so if $X$ is a complete metric space under one of the two uniformly equivalent metrics, it is complete under the other. Any two compatible metrics on a compact space are easily verified to be uniformly equivalent.
{crllr}
Let $(X,d)$ be a complete separable metric space of uncountable cardinality. There exists a metric $\rho$ on $X$ uniformly equivalent to $d$ and such that the $k$-NN classifier is not universally consistent in the (complete, separable) metric space $(X,\rho)$, under any choice of sequence $k_{n}\to\infty$, $k_{n}/n\to 0$.
In particular, already the following feels a bit counter-intuitive.
{crllr}
There is a metric $\rho$ on the real line $\mathbb{R}$}, uniformly equivalent to the usual metric, and such that the $k$-NN classifier is not universally consistent in the (complete, separable) metric space $({\mathbb{R}},\rho)$, under any choice of sequence $k_{n}\to\infty$, $k_{n}/n\to 0$.
We will first collect some results used in the proof.
{xmpl}
[Davies [10], Theorem II] There exists a compact metric space $\Omega$ of diameter one and two Borel measures $\mu_{1},\mu_{2}$ with $\mu_{1}(\Omega)=1/3$, $\mu_{2}(\Omega)=2/3$, such that $\mu_{1}(B)=\mu_{2}(B)$ for every closed ball in $\Omega$ of radius less than one.
The distances in the space $\Omega$ all take values of the form $2^{-i}$ (ibid., p. 159-8), which implies that $\Omega$ is zero-dimensional and, as by the construction it has no isolated points, is homeomorphic to the Cantor set $\{0,1\}^{{\mathbb{N}}}$ by the Brouwer theorem ([17], Theorem 7.4).
Denote $\eta_{1},\eta_{2}$ the Radon-Nikodým derivatives of $\mu_{1}$ and $\mu_{2}$, respecively, with regard to the probability measure $\mu_{1}+\mu_{2}$. They are obviously different between themselves, so at least one of $\eta_{1},\eta_{2}$ fails the Lebesgue differentiation property.
The theorem of Preiss–Assouad–Quentin de Gromard (see Subsection 2.2) implies that the Davies metric space $\Omega$ is not sigma-finite dimensional.
The following is a classical result from the early days of general topology and descriptive set theory.
{thrm}
[See e.g. [17], Theorems 6.2 and 6.4] Every complete separable metric space of uncountable cardinality contains a homeomorphic copy of the Cantor space.
And here is one of the recent variations on the theorem of Hausdorff about extending metrics from subspaces (see e.g. [32]).
{thrm}
[Nguyen To Nhu, [31], Corollary 3.5] Let $K$ be a compact subspace of a metric space $(X,d)$, and let $\rho$ be a compatible metric on $K$. Then $\rho$ extends to a metric $\bar{\rho}$ on $X$ uniformly equivalent to $d$.
Proof of Corollary 6.
Let $X=(X,d)$ be a complete separable metric space of uncountable cardinality. Identify the Davies space $\Omega\cong\{0,1\}^{{\mathbb{N}}}$ (Example 6) with a topological subspace of $X$ (Theorem 6), and extend the metric from $\Omega$ to a metric $\rho$ over all of $X$ that is uniformly equivalent to $d$ (Theorem 6). Since the metric on $\Omega$ is not sigma-finite dimensional, the same is true for its extension $\rho$. According to our Theorem 1, the $k$-NN classifier is not universally consistent in $(X,\rho)$. ∎
{rmrk}
Our Corollary 6 was inspired by the (apparently unpublished) result of Andretta, Camerlo and Costantini ([2], slide 12): every separable complete metric space equipped with a non-atomic probability measure, $(X,d,\mu)$, admits a complete compatible metric $\rho$ such that $(X,\rho,\mu)$ fails the Lebesgue–Besicovitch differentiation theorem.
The result uses a strenthened form of Theorem 6, namely, Theorem 3.1 in [1], affirming that the Cantor set with the canonical Borel probability measure admits a measure-preserving homeomorphic embedding into any complete separable metric space $X$ equipped with a non-atomic measure $\mu$ if for some Borel set, $A\subseteq X$, one has $1<\mu(A)<\infty$. Another ingredient is the example of Käenmäki, Rajala and Suomala from [16], mentioned above as Example 5, where a metric with a doubling property is constructed on the Cantor set with regard to which the standard Borel probability measure fails the Lebesgue–Besicovitch property. Here the doubling property is not used, so the (much earlier) Davies example could have been used instead.
Let us remark that if instead of the Hausdorff extension theorem one applies the result of Nguyen To Nhu [31] (Theorem 6), one can further strengthen the above result by Andretta, Camerlo and Costantini by assuming the metric $\rho$ to be uniformly equivalent to $d$ as well.
Of course all we needed in order to apply our Theorem 1, is the existence on $X$ of a uniformly equivalent metric $\rho$ that is not sigma-finite dimensional, so a much simpler argument suffices.
7. The $1$-nearest neighbour classifier
Here we aim to prove Theorem 1. Let us recall that the result states the equivalence, for every complete metric space $\Omega$, of the following two conditions between themselves and with all the three equivalent conditions in Theorem 1:
  1. (4)
    Under every distribution of the labelled data on $\Omega\times\{0,1\}$, the misclassification error of the $1$-nearest neighbour rule, asymptotically as the size of an i.i.d. labelled sample goes to infinity, does not exceed $2\ell^{\ast}$, where $\ell^{\ast}$ is the Bayes error.
  2. (4)
    Under every distribution of the labelled data on $\Omega\times\{0,1\}$ whose Bayes error is zero (that is, the regression function is deterministic), the $1$-nearest neighbour rule is consistent when the size of an i.i.d. labelled sample goes to infinity.
[…]
Let $\Omega$ be a complete separable metric space, equipped with a Borel probability measure $\mu$ and a regression function $\eta$ that satisfied the strong Lebesgue–Besicovitch differentiation property with regard to $\mu$. Let $(X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n})$ be a random i.i.d. $n$-sample, where $X_{i}\in\Omega$ are i.i.d. random variables following the law $\mu$ and $Y_{i}$ are independent Bernoulli random variables with probabilities of success $\eta(X_{i})$. To break the distance ties, we use auxiliary i.i.d. real random variables $\xi_{1},\ldots,\xi_{n}$ following some non-atomic law and independent of the remaining random variables.
A way to explicitly generate $Y_{i}$ is to fix a sequence of i.i.d. random variables $Z_{i}$ uniformly distributed on the interval $[0,1]$, independent of $X_{i}$ and of $\xi_{i}$, and to set
$Y_{i}={\mathbb{I}}_{Z_{i}\leq\eta(X_{i})}.$
Now fix $x\in\Omega$ and define auxiliary i.i.d. Bernoulli random variables
$Y_{i}^{\prime}={\mathbb{I}}_{Z_{i}\leq\eta(x)},$
all having the same probability of success $\eta(x)$. They are independent of $X_{1},\ldots,X_{n}$ and of $\xi_{1},\ldots,\xi_{n}$, and are correlated with the labels $Y_{1},Y_{2},\ldots,Y_{n}$.
Let $X_{(1)}(x)$ denote the nearest neighbour to $x$ within the sample $(X_{1},\ldots,X_{n})$ chosen, if need be, with the help of the auxiliary variables $(\xi_{i})$. Denote $Y_{(1)}(x)$ and $Y^{\prime}_{(1)}(x)$ the random labels as above assigned to $X_{(1)}(x)$. In particular, $Y_{(1)}(x)$ is the predictor made by the $1$-NN classifier for the point $x$ on the random input $(X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n})$.
The following is Lemma 5.2 in [13] in the particular case $k=1$. (It was stated for the Euclidean space, but the nature of the distance in the domain does not matter.)
{lmm}
${\mathbb{E}}\left|Y_{(1)}(x)-Y^{\prime}_{(1)}(x)\right|\leq{\mathbb{E}}\left|\eta(x)-\eta(X_{(1)}(x))\right|$.
Proof.
Conditioning on the sample $X_{1},\ldots,X_{n},\xi_{1},\ldots,\xi_{n}$, we end up with two Bernoulli random variables of the form $Y^{\prime}={\mathbb{I}}_{Z\leq p}$ and $Y={\mathbb{I}}_{Z\leq p_{1}}$ and just need to upper bound ${\mathbb{E}}\lvert Y-Y^{\prime}\rvert=P[Y\neq Y^{\prime}]$ by $\lvert p-p_{1}\rvert$. ∎
{lmm}
As $n\to\infty$, ${\mathbb{E}}\left|\eta(X)-\eta(X_{(1)}(X))\right|\to 0$.
Proof.
This is the second part of the proof of Theorem 2.2 in [7] starting with p. 344-3, again corresponding to the particular case of fixed $k=1$. The assumption $k\to\infty$ is essential in the first part of the proof, but in the second part, only the assumption $k/n\to 0$ is needed to invoke the Cover–Hart lemma at the very end, subject to a slight adjustment in the proof as indicated to us by Frédéric Cérou (see Subsection 2.3). ∎
Let now $(X,Y)$ be a random pair following the same law as $(X_{1},Y_{1})$ and independent of the rest of random variables. Define the misclassification error of the $1$-NN classifier
$\ell_{n}=P[Y_{(1)}(X)\neq Y],$
and introduce the misclassification error of the auxiliary classifier making for the point $x$ the prediction $Y^{\prime}_{(1)}(x)$:
$\ell^{\prime}_{n}=P[Y^{\prime}_{(1)}(X)\neq Y].$
{lmm}
[Cf. top of the p. 70 in [13] in the Euclidean case] ${\mathbb{E}}\left|\ell_{n}-\ell^{\prime}_{n}\right|\to 0$ when $n\to\infty$.
Proof.
Using Lemmas 7 and 7,
${\mathbb{E}}\left|\ell_{n}-\ell^{\prime}_{n}\right|$ $\leq P[Y^{\prime}_{(1)}(x)\neq Y_{(1)}(X)]$
$\leq{\mathbb{E}}\left|\eta(X)-\eta(X_{(1)}(X))\right|$
$\to 0.$
It remains only to estimate the error of the auxiliary classification rule $Y^{\prime}_{(1)}(X)$, which is straightforward, as now we deal with the probability that two independent identically distributed Bernoulli variables take different values:
$\ell^{\prime}_{n}$ $=P[Y^{\prime}_{(1)}(X)\neq Y]$
$={\mathbb{E}}\left(2\eta(X)(1-\eta(X))\right)$
$\leq 2{\mathbb{E}}\min\{\eta(X),1-\eta(X)\}$
$=2\ell^{\ast}.$
(Cf. Theorem 5.1 in [13].)
Thus, asymptotically, the classification error of the $1$-NN rule in a complete separable metric space satisfying any one of the equivalent conditions in Theorem 1 (for instance, being sigma-finite dimensional in the sense of Nagata if you will) is bounded by twice the Bayes error of the labelled data distribution:
$\ell^{\ast}(\mu,\eta)\leq\limsup_{n\to\infty}\ell_{n}\leq 2\ell^{\ast}(\mu,\eta).$